### v1

@@ 5.1 Synthetic 3 Variables
Using more flexible estimators may introduce some probability of error (section 2.4, appendix A). First, we answer the question: if the data generating process matches our model, how informative is the MAP value (eq. 1) about the true causal structure? (appendix A). To estimate the error only due to overlap in posteriors of different causal models, we perform discrete model comparison (labelled DGP-CDE, see section C.1 for more details). This involves enumerating and computing the posterior of every causal model. We study the 3 variable case, where it is possible to enumerate every possible causal structure, 25 in total. Here, data is generated from our GP-CDE model, and from each of the 6 distinct causal structures for 3 variables (up to permutations of the variables, details in appendix E). We generate 5 datasets of 1000 samples for each of the graphs leading to 30 datasets in total.
@@
The results for the 3 variable case can be seen in fig. 2 (a). Here, we can see that the discrete comparison DGP-CDE achieves very good performance. This shows that in the GP-CDE model, the overlap between the posteriors of different causal structures is low and that our inference does not contribute greatly to errors. Thus, if the GP-CDE model is a good descriptor of the dataset at hand, we can expect the MAP value to identify the true causal direction with high probability. In Figure 2 (a) the CGP-CDE performs worse across all metrics than the DGP-CDE. This shows that we can expect a higher error due to optimisation of the continuous relaxation in the CGP-CDE.
5.2 Synthetic 10 Variables
@@
Next, we analyse the performance on a higher number of variables. Here, performing discrete model comparison is too costly as the number of DAGs is prohibitively large, so we use our continuous relaxation, labelled CGP-CDE. We generate 10 node graphs using Erdos-Renyi (ER) and scale-free (SF) (Barabási and Albert, 1999) schemes with 15 expected edges. In ER graphs, each edge is generated independently while SF graphs promote a few nodes with a high number of edges. The data is generated from randomly initialised neural networks with noise included as an input. This ensures that the data is not sampled from any of the baselines or the CGP-CDE model. Performing well in such cases showcases the advantage of the added flexibility in the Bayesian approach. Data generation details are in appendix E.
The results for the 10 variable synthetic datasets can be seen in fig. 2 (b) and (c). Compared to the DGP-CDE results of section 5.1, we can expect a higher number of errors due to the continuous relaxation, as well as the shift in the data generation process. Nevertheless, for the ER graph, the CGP-CDE method outperforms the baselines in terms of the mean SHD, SID and F1 scores. For the SF graphs, CGP-CDE performs competitively, achieving better mean SID and F1 scores for most methods except for NOGAM and SCORE. In the SF graphs, a few variables cause the rest of the variables. We hypothesise that this structure benefits the topological search that NOGAM and SCORE perform.
5.3 Syntren and Sachs
Table 1: Results for Sachs dataset (Sachs et al., 2005). Lower is better for SHD, SID, higher is better for F1.
Method SHD $\downarrow$ SID $\downarrow$ F1 $\uparrow$
NOTEARS $\mathbf{14}$ $\mathbf{49}$ $0.36$
GraN-DAG $19$ $57$ $0.30$
GraN-DAG ++ $18$ $55$ $0.31$
CAM $18$ $56$ $0.36$
SCORE $23$ $59$ $0.15$
NOGAM $23$ $59$ $0.15$
CGP-CDE $17$ $59$ $\mathbf{0.48}$
Syntren is a semi-synthetic dataset that takes real biological graphs and generates datasets according to a simulator. There are 10 datasets of 20 nodes with 500 samples each. The Sachs dataset (Sachs et al., 2005) is a real dataset that measures the expression levels of proteins. It contains $853$ observational samples and 11 nodes. Details for both are in appendix E.
The results on Syntren can be seen in fig. 2 (d). Here, the CGP-CDE performs competitively in terms of F1 and SHD scores, only losing out to GraN-DAG and GraN-DAG++ in SID. However, GraN-DAG and GraN-DAG++ perform worse in terms of SHD and F1. The SID score does not penalise extra edges. Syntren has on average $23.5$ edges. While CGP-CDE predicts $45.4$, GraN-DAG and GraN-DAG++ predict $79.7$ and $74.8$ respectively. The results for Sachs can be seen in table 1. The CGP-CDE outperforms all the methods here in terms F1 scores. NOTEARS performs well on SHD and SID but only predicts 4 edges, whereas CGP-CDE predicts 15. The ground truth has 18 edges.
@@ 6 Conclusion
We have introduced a continuous relaxation approach to Bayesian model selection for multivariate causal discovery. This method casts the problem as a single differentiable optimisation task, allowing the use of more flexible models compared to traditional methods. Our approach demonstrates competitive performance across various synthetic and real-world datasets, highlighting its versatility. By allowing scalable causal discovery without restrictive assumptions, we illustrate that Bayesian model selection is effective and could allow for more practical causal discovery.
Appendix A Discussion on Identifiability and Probability of Error
With only observational data it is possible to recover the Markov equivalence class (MEC) of a causal structure with only the Markov and faithfulness assumptions [Pearl, 2009]. To recover a DAG within an MEC, additional assumptions are required. Here we provide background on the assumptions that provide strict identifiability and the Bayesian model selection approach that allows for more flexible model classes but may incur some probability of error.
The probability of error can be defined as the error made by a model and a decision rule when multiple datasets are sampled from the model itself [Guyon et al., 2019, Dhir et al., 2024]. Restrictive assumptions on models such as additive-noise (ANM), provide a strict notion of identifiability with maximum likelihood [Zhang et al., 2015], where the probability of error is 0 (fig. 3 (b)). This suggests, that if the data is sampled from an ANM, the causal direction can be recovered with probability 1 (in the population setting). Other examples of restricted model classes include post non-linear noise models (PNL) [Zhang and Hyvarinen, 2012], and location-scale noise models [Immer et al., 2023]. However, if the data generating process deviates from these restrictions, then the above guarantee no longer holds.
Bayesian model selection preserves the guarantees of maximum likelihood based approaches if suitably restricted model classes are used [Dhir et al., 2024, Prop. 4.5]. However, for more flexible model classes, maximum likelihood based approaches are completely non-informative within a Markov equivalence class (fig. 3(a)) [Zhang et al., 2015, Lemma 1]. For these flexible model classes, Bayesian model selection can still be informative [Dhir et al., 2024, Prop. 4.7]. Bayesian model selection involves placing priors on a model’s functions or parameters, thereby assigning priors over its distributions. These priors are constructed to encode the Independent Causal Mechanisms (ICM) assumption in a specific causal direction; changing the causal direction alters how the ICM assumption is applied. The marginal likelihood then estimates the probability that the observed data were generated under the model’s prior and corresponding ICM assumption. Because the marginal likelihood must sum to one over all possible datasets (see fig. 3 (c)), it cannot assign equal probability to different causal directions under the same prior and its ICM assumption (unless under certain conditions hold [Dhir et al., 2024, Corollary. 4.8]). Here, the probability of error may no longer be 0 as the posteriors may overlap(fig. 3(c)). The advantage of the Bayesian approach is that the more flexible assumptions may mitigate misspecification in cases where more flexibility is required to model the data well and the guarantees of restricted models do not hold anyway. This probability of error can be estimated by sampling datasets from the model and classifying the causal direction [Dhir et al., 2024, Sec. 4.3]. We perform this in section 5.1.
$\mathcal{D}$$\max_{\theta,\ \phi}\ p(\mathcal{D}|\theta,\phi,\mathcal{M}_{X\rightarrow Y})$$\max_{\theta,\ \phi}\ p(\mathcal{D}|\theta,\phi,\mathcal{M}_{Y\rightarrow X})$
$\mathcal{D}$$\max_{\theta,\ \phi}\ p(\mathcal{D}|\theta,\phi,\mathcal{M}^{\prime}_{X\rightarrow Y})$$\max_{\theta,\ \phi}\ p(\mathcal{D}|\theta,\phi,\mathcal{M}^{\prime}_{Y\rightarrow X})$
$\mathcal{D}$$p(\displaystyle\mathcal{D}|\mathcal{M}_{X\ \rightarrow Y})$$p(\displaystyle\mathcal{D}|\mathcal{M}_{Y\ \rightarrow X})$
Figure 3: Reproduced from Dhir et al. [2024] with permission from authors. (a) Maximum likelihood based methods cannot distinguish causal models that are within a Markov Equivalence Class. (b) Making suitable restrictions allows for maximum likelihood to identify the causal direction, but restricts the datasets that can be modelled. (c) Bayesian model selection relaxes these restrictions by incorporating priors that encode the Independent Causal Mechanisms (ICM) assumption. This approach offers greater modelling flexibility and enables estimation of how likely it is that a dataset was generated under the prior and encoded ICM assumption.
Appendix B CGP-CDE Details

### v2

@@ 6.1 Synthetic 3 Variables
Using more flexible estimators may introduce some probability of error (Section 3). In the two variable case, Dhir et al. (2024) showed that the probability of error for our model class is low; we show that this is true for the multivariate case as well. To estimate the error only due to overlap in posteriors of different causal models, we perform discrete model comparison (labelled DGP-CDE, see Section D.1 for more details). This involves enumerating and computing the posterior of every causal model. We study the 3 variable case, where it is possible to enumerate every possible causal structure, 25 in total. Here, as required by the probability of error, data is generated from our GP-CDE model (details in Appendix F). We generate five datasets of 1000 samples for each of the six graphs that are unique up to a permutation of the variables.
@@
The results for the 3 variable case can be seen in Figure 1. Here, we can see that the discrete comparison DGP-CDE achieves very good performance. This shows that in the GP-CDE model, the probability of error in the multivariable case is low and that our inference does not contribute greatly to errors. Thus, if the GP-CDE model is a good descriptor of the dataset at hand, we can expect the MAP value to identify the true causal direction with high probability. We also compare our continuous relaxation model against the discrete model that enumerates each graph type. In Figure 1, the performance of CGP-CDE shows that we can expect a higher error due to optimisation of the continuous relaxation in the CGP-CDE. However, the CGP-CDE still performs competitively compared to other baselines.
6.2 Synthetic Higher Variable
@@
Next, we analyse the performance on a higher number of variables with varying graph density. Here, performing discrete model comparison is too costly as the number of DAGs is prohibitively large, so we use our continuous relaxation, labelled CGP-CDE. We generate five datasets for each graph size 50 using Erdos-Rényi (ER) and scale-free (SF) (Barabási & Albert, 1999) graph generation schemes, with expected total edge counts of 1D and 4D, respectively. In ER graphs, each edge is generated independently, while SF graphs promote a few nodes with a high number of edges. To test our claim that Bayesian model selection allows for more flexible models, we ensure that the data is not generated from our model. The data is generated from randomly initialised neural networks with noise included as an input. Performing well in such cases showcases the advantage of the added flexibility in the Bayesian approach. Data generation details are in Appendix F. To show that our method also performs well with a lower number of variables and samples, we show results for a graph size of $20$ in Section I.3 and Section I.2. We also perform an ablation study with ANM datasets in Section I.1.
The full results are in Appendix I, and we show the 50 variable ER4 results in Figure 2(a). Compared to the DGP-CDE results of Section 6.1, we can expect a higher number of errors due to the continuous relaxation and the shift in the data generation process. Nevertheless, the CGP-CDE outperforms all other baselines, especially in terms of SHD and F1 scores, and the increase in performance compared to the baselines is greater for these larger graphs. While both SDCD and CGP-CDE employ flexible function approximators and the same acyclic regulariser, the superior performance of CGP-CDE suggests that Bayesian methods are more effective at distinguishing causal structures. In contrast, the poor performance of DiBS (MAP estimation) can be attributed to poor inference in Bayesian neural networks.
6.3 Syntren
Syntren is a gene-regulatory network simulator that generates gene expression data from real biological graphs. There are $10$ datasets of 20 nodes with 500 samples each. The results on Syntren can be seen in Figure 2(b). Here, the CGP-CDE outperforms other methods, especially in terms of the SID and F1 scores.
@@ 7 Conclusion
In this work, we have shown why Bayesian model selection works, and that it is a highly effective approach for causal discovery in multivariable settings. Then, we propose a scalable model that outperforms existing Bayesian and non-Bayesian methods across multiple settings. The consistent performance of our approach highlights the advantages of framing causal discovery as a Bayesian model selection problem. We believe this perspective addresses a critical limitation in causal discovery—namely, the inability to use flexible models, paving the way for more practical and widely applicable causal inference in real-world scenarios.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Acknowledgements
R.S. was supported by the Wellcome Trust [222836/Z/21/Z]. B.G. acknowledges the support of the UKRI AI programme, and the Engineering and Physical Sciences Research Council, for CHAI - EPSRC Causality in Healthcare AI Hub (grant no. EP/Y028856/1). A.K. is supported by UKRI (grant number EP/S023356/1), as part of the UKRI Centre for Doctoral Training in Safe and Trusted AI. A.D. acknowledges the support of G-Research’s March 2025 grant.
Appendix A Additional Background
A.1 Causal Model Basics
The causal model satisfies the Independent Causal Mechanism (ICM) assumption by construction (Janzing & Schölkopf, 2010). Given a graph $\mathcal{G}$, this means that changing a term in the causal factorisation of the joint $\prod_{i=1}^{D}P_{i}(X_{i}|X_{\mathrm{PA}_{\mathcal{G}}(i)})$ does not change any of the other terms. This is achieved by assuming that the choice of conditional distributions for each variable $X_{i}$ from $\mathcal{C}_{i|\mathrm{PA}_{\mathcal{G}}(i)}$ is independent of the choice for other variables $X_{j}$ from $\mathcal{C}_{j|\mathrm{PA}_{\mathcal{G}}(j)}$ with $j\neq i$. This property is assumed to hold in the causal factorisation but does not necessarily hold in any other factorisation of the joint (Peters et al., 2017). This can be seen in the two variable system with graph $X\to Y$. If the terms in $P(X)P(Y|X)$ are chosen independently, the choice of the terms in $P(Y)P(X|Y)$ are tied to each other. Changing $P(X)$ or $P(Y|X)$ in general will require changes in both $P(Y)P(X|Y)$ to maintain equality.
As variables are dependent only on their parents in a graph by construction, a causal model satisfies the Markov assumption. This states that d-separations in $\mathcal{G}$ between variables implies conditional independence between the same variables (Pearl, 2009). We also assume faithfulness: any conditional independences between variables are represented by d-separations in $\mathcal{G}$ (Pearl, 2009). Together, faithfulness and Markov assumptions imply a one-to-one relationship between independences in the distributions and d-separations in the graph (Peters et al., 2017). These assumptions are enough to recover a causal structure from data up to a Markov equivalence class (MEC), that is, the class of causal structures that have the same d-separations.
A.2 Continuous Optimisation for Learning Causal Structure
That is, finding the graph that optimises the score under the constraint that the graph must be a valid DAG. This converts causal structure learning into a single optimisation problem, tackling the scalability issue. However, ensuring that the graph is acyclic is non-trivial with gradient based methods. Zheng et al. (2018) first solve this by encoding a graph as a weighted adjacency matrix $\mathbf{A}\in\mathbb{R}^{D\times D}_{\geq 0}$ and using a measure of acyclicity $h(\mathbf{A})$ as a constraint. The function $h:\mathbf{A}\to\mathbb{R}_{\geq 0}$ is a measure of the number of weighted directed walks that allow for returning to a starting node. Hence, $h(\mathbf{A})=0$ implies that there are no directed cycles in the adjacency $\mathbf{A}$ and that the graph implied by $\mathbf{A}$ is a valid DAG.
Numerous variations on this approach have been proposed to improve the scalability and stability of optimisation (Lee et al., 2019; Bello et al., 2022; Wei et al., 2020; Nazaret et al., 2024). We use the spectral acyclicity constraint method in Equation 12, which is equivalent to the largest eigenvalue magnitude of $\mathbf{A}$, due to its stability and scalability (Lee et al., 2019; Nazaret et al., 2024). This constraint uses the fact that $\mathbf{A}$ is acyclic if and only if all its eigenvalues are zero (Cvetković et al., 1980).
The gradients of the spectral acyclicity constraint can be easily calculated as
$\nabla h(\mathbf{A})=\frac{v_{d}u_{d}^{T}}{v_{d}^{T}u_{d}},$ (15)
where $u_{d}$ and $v_{d}$ are the right and left eigenvectors associated with $\lambda_{d}(\mathbf{A})$ (Magnus, 1985). We follow the example of Nazaret et al. (2024) and use the power iteration method to estimate these eigenvectors in $\mathcal{O}(d^{2})$ rather than finding them exactly, which costs $\mathcal{O}(d^{3})$.
Equation 14 can now be written as an optimisation problem with the differentiable constraint $h(\mathbf{A})=0$. Nazaret et al. (2024) use a penalty method, writing the objective as:
$\operatorname*{\arg\!\max}_{\mathbf{A}}\mathcal{S}(\mathcal{G}^{\mathbf{A}})-\gamma_{t}h(\mathbf{A}),$ (16)
where the weighting $\gamma_{t}$ of the penalty $h(\mathbf{A})$ is increased by $\rho$ each epoch. A solution $\mathbf{A}^{*}$ is found when $h(\mathbf{A}^{*})<\tau$ for some convergence threshold $\tau>0$.
Appendix B Proof of Distinguishability with Multiple Variables
Bayesian model selection uses the marginal likelihood (equivalent to the posterior over models) to distinguish between Bayesian causal models with different causal graphs. Thus, following Dhir et al. (2024), we will provide conditions such that two Bayesian causal models will have different distributions over marginal likelihood values. This implies that there will always exist datasets such that the marginal likelihood values of two Bayesian causal models differ. We will do this by defining two Bayesian causal models as Bayesian distribution-equivalent if they have the same distribution over marginal likelihood values, and then provide conditions such that two Bayesian causal models cannot be Bayesian distribution-equivalent. Note that this can be viewed as a version of observation equivalence (Pearl, 2009), but for Bayesian causal models. If models are not observationally equivalent, they can be learnt from observational data (for example, by using methods such as PC (Spirtes et al., 2000)). A similar reasoning holds for Bayesian distribution equivalence.
We split the proof into two parts, first, we will show that Bayesian model selection can effectively identify up to a Markov equivalence class. This is known for certain classes of models, for example, curved exponential models (Chickering, 2002), but it has not been shown for the general non-parametric case. Second, we will provide conditions such that two Bayesian causal models within the same Markov equivalence class are not Bayesian distribution-equivalent. Here, there may be some probability of error depending on the overlap of the posteriors that can be computed empirically. This second result is implied by the result in Dhir et al. (2024, Proposition 4.7).
We define Bayesian distribution-equivalent models. These are the Bayesian counterparts of models that are unidentifiable. That is, given any dataset, these models will achieve the same marginal likelihood score.
Definition B.1.
(Dhir et al., 2024, Definition 4.4) Given two Bayesian causal models $(\mathcal{M}_{\mathcal{G}},\pi_{\mathcal{G}})$, $(\mathcal{M}_{\mathcal{H}},\pi_{\mathcal{H}})$, say they are Bayesian distribution-equivalent if $P\mathopen{}\mathclose{{}\left(\cdot\mid\mathcal{M}_{\mathcal{G}}}\right)=P\mathopen{}\mathclose{{}\left(\cdot\mid\mathcal{M}_{\mathcal{H}}}\right)$, i.e. for all $N\in\mathbf{N}$, and for all $\mathopen{}\mathclose{{}\left(\mathbf{x}^{N},\mathbf{y}^{N}}\right)\in\mathopen{}\mathclose{{}\left(\mathcal{X}\times\mathcal{Y}}\right)^{N}$, it holds that $p\mathopen{}\mathclose{{}\left(\mathbf{x}^{N},\mathbf{y}^{N}\mid\mathcal{M}_{\mathcal{G}}}\right)=p\mathopen{}\mathclose{{}\left(\mathbf{x}^{N},\mathbf{y}^{N}\mid\mathcal{M}_{\mathcal{H}}}\right)$.
We proceed by showing that two Bayesian causal models with graphs in different MECs cannot be Bayesian distribution equivalent. Further, if the estimate of the Bayes factor (Kass & Raftery, 1995) (ratios of marginal likelihood) is consistent, it can identify the correct graph.
Lemma B.2.
Assume faithfulness and two Bayesian causal models $(\mathcal{M}_{\mathcal{G}},\pi_{\mathcal{G}})$, and $(\mathcal{M}_{\mathcal{H}},\pi_{\mathcal{H}})$. If $\mathcal{G}$ and $\mathcal{H}$ are in separate Markov equivalence classes (MEC), then $\mathcal{F}_{\mathcal{G}}\cap\mathcal{F}_{\mathcal{H}}=\emptyset$, where $\mathcal{F}_{\mathcal{G}}$ and $\mathcal{F}_{\mathcal{H}}$ are the sets of all joints expressible by $\mathcal{M}_{\mathcal{G}}$ and $\mathcal{M}_{\mathcal{H}}$ respectively.
Proof.
The Markov and Faithfulness assumptions together means that $X_{i}\perp\!\!\!\!\perp_{P}X_{j}|X_{k}\iff X_{i}\perp\!\!\!\!\perp_{\mathcal{G}}X_{j}|X_{k}$, where $\perp\!\!\!\!\perp_{P}$ denotes distributional independence, and $\perp\!\!\!\!\perp_{\mathcal{G}}$ signifies d-separation in graph $\mathcal{G}$. Hence, as $\mathcal{G}$ and $\mathcal{H}$ are in separate MECs, there exists some $X_{i}\perp\!\!\!\!\perp_{\mathcal{G}}X_{j}|X_{k}$ such that $X_{i}\not\perp\!\!\!\!\perp_{\mathcal{H}}X_{j}|X_{k}$. This implies that for every $P\in\mathcal{F}_{\mathcal{G}}$, we have that $X_{i}\perp\!\!\!\!\perp_{P}X_{j}|X_{k}$, and for every $Q\in\mathcal{F}_{\mathcal{H}}$ we have that $X_{i}\not\perp\!\!\!\!\perp_{Q}X_{j}|X_{k}$ Hence $P\notin\mathcal{F}_{\mathcal{H}}$. ∎
Theorem B.3.
Assume faithfulness and two Bayesian causal models $(\mathcal{M}_{\mathcal{G}},\pi_{\mathcal{G}})$, and $(\mathcal{M}_{\mathcal{H}},\pi_{\mathcal{H}})$. If $\mathcal{G}$ and $\mathcal{H}$ are in separate Markov equivalence classes (MEC), then they are not Bayesian distribution equivalent for any choices of priors $\pi_{\mathcal{G}},\pi_{\mathcal{H}}$. Assume $N$ samples are generated from one of the models. If the estimate $\frac{P(X^{N}|\mathcal{M}_{\mathcal{G}})}{P(X^{N}|\mathcal{M}_{\mathcal{H}})}$ is consistent as $N\to\infty$, the probability of error between these two models is 00.
Proof.
The first claim follows directly from Lemma B.2 and Dhir et al. (2024, Proposition 4.5). Given that $\mathcal{F}_{\mathcal{G}}\cap\mathcal{F}_{\mathcal{H}}=\emptyset$, only one of the models can be the true model. Consistency implies $\frac{P(X^{N}|\mathcal{M}_{\mathcal{G}})}{P(X^{N}|\mathcal{M}_{\mathcal{H}})}\to\infty$ if the data generating distribution $P\in\mathcal{F}_{\mathcal{G}}$, and $\frac{P(X^{N}|\mathcal{M}_{\mathcal{G}})}{P(X^{N}|\mathcal{M}_{\mathcal{H}})}\to 0$ if the data generating distribution $P\in\mathcal{F}_{\mathcal{H}}$ (Walker, 2004). Hence, the marginal likelihood is higher for the correct model for each distribution in the union of $\mathcal{F}_{\mathcal{G}}\cup\mathcal{F}_{\mathcal{H}}$ leading to a probability of error of 00. ∎
Consistency in the above requires certain conditions on the prior (Walker et al., 2004, Theorem 1). Suppose the data is generated from $X^{N}\sim P_{0}$, we require that for the correct model, suppose it is $\mathcal{M}_{\mathcal{G}}$, the prior puts non-negligible mass around $P_{0}$: $\pi_{\mathcal{G}}(\{P:\text{KL}[P_{0}\|P]<\epsilon\})>0$ for all $\epsilon>0$ (Walker et al., 2004). Second, the posterior for both models cannot paradoxically concentrate in a region of negligible prior mass: $\lim\inf_{N}\text{KL}[P_{0}\|P_{NA}]\geq\epsilon$ for all $\epsilon>0$ where $A:=\{P:\text{KL}[P_{0}\|P]>\epsilon\}$. $P_{NA}$ is the posterior (over $N$ samples) computed from the prior but restricted to the set of densities $A$, which are $\epsilon$ far from the true density $P_{0}$ as measured by the KL-divergence (Walker, 2004). This condition stops the wrong model (which has negligible prior mass around the true density) from concentrating around the true density for a large enough sample size.
The condition for two causal models that $\mathcal{F}_{\mathcal{G}}\cap\mathcal{F}_{\mathcal{H}}=\emptyset$ also holds when the causal models are identifiable (Guyon et al., 2019, Ch. 2). Hence, Bayesian model selection can recover the causal structure if the causal models are identifiable.
Corollary B.4.
Assume two Bayesian causal models $(\mathcal{M}_{\mathcal{G}},\pi_{\mathcal{G}})$, and $(\mathcal{M}_{\mathcal{H}},\pi_{\mathcal{H}})$ where the underlying causal models are identifiable: $\mathcal{F}_{\mathcal{G}}\cap\mathcal{F}_{\mathcal{H}}=\emptyset$. If the estimate $\frac{P(X^{N}|\mathcal{M}_{\mathcal{G}})}{P(X^{N}|\mathcal{M}_{\mathcal{H}})}$ is consistent as $N\to\infty$, the probability of error between these two models is 00.
Next, we consider the case of two Bayesian causal models where the two graphs are in the same MEC. Without restrictions, it is well known that in this case there may exist distributions in two Bayesian causal models that can express a given dataset (Pearl, 2009). Thus, the result from Lemma B.2 does not hold. The key idea that allows for distinction here is that the ICM assumption (as represented by factorised priors) may only hold in the causal factorisation of a Bayesian causal model. Further, the distributions of marginal likelihood values are sensitive to the ICM assumption. As the causal factorisation of the two Bayesian causal models differs, they will have a different distribution over marginal likelihood values.
We first formalise the ICM assumption as the separability of the priors with respect to a certain factorisation of the joint.
Definition B.5.
(Dhir et al., 2024, Definition 4.2) Given a Bayesian causal model $(\mathcal{M}_{\mathcal{G}},\pi_{\mathcal{G}})$, a prior $\pi$ is separable with respect to $\mathcal{C}_{\mathcal{G}}$ if it factorises $\prod_{i\in\mathcal{V}}\pi_{i}$ such that $\pi_{i}\in\mathcal{P}(\mathcal{C}_{i|\mathrm{PA}_{\mathcal{G}}(i)})$, where $\mathcal{P}$ is the set of all distributions over an object.
That is, if $\pi$ is separable with respect to $\mathcal{C}_{\mathcal{G}}$, the joint can be written as $\prod_{i\in\mathcal{V}}P_{i}(X_{i}|X_{\mathrm{PA}_{\mathcal{G}}(i)})\pi(\mathrm{d}P_{i})$, with $P_{i}\in\mathcal{C}_{i|\mathrm{PA}_{\mathcal{G}}(i)}$. Separability of the prior (and hence the ICM condition) may hold in multiple factorisations. This is the case with normalised linear Gaussian models (Dhir et al., 2024, Appendix D.2). To formalise this, we reiterate the notion of separability-compatibility.
Definition B.6.
(Dhir et al., 2024, Definition 4.6) Given two Bayesian causal models $(\mathcal{M}_{\mathcal{G}},\pi_{\mathcal{G}})$, and $(\mathcal{M}_{\mathcal{H}},\pi_{\mathcal{H}})$. Denote $\gamma:\mathcal{C}_{\mathcal{G}}\to\mathcal{C}_{\mathcal{H}}$ a bijection such that for any $P\in\mathcal{C}_{\mathcal{G}}$ such that $Q:=\delta(P)\in\mathcal{C}_{\mathcal{H}}$, there holds an equality of joint measures $\prod_{i\in\mathcal{V}}P_{i}(X_{i}|X_{\mathrm{PA}_{\mathcal{G}}(i)})=\prod_{i\in\mathcal{V}}Q_{i}(X_{i}|X_{\mathrm{PA}_{\mathcal{H}}(i)})$. The two Bayesian causal models are separable-compatible if: i) the pushforward $\pi_{\mathcal{G}}\circ\gamma^{-1}$ is separable with respect to $\mathcal{C}_{\mathcal{H}}$, ii) $\pi_{\mathcal{H}}\circ\gamma$ is separable with respect to $\mathcal{C}_{\mathcal{G}}$.
The operation $\delta$ is simply the transformation by Bayes’ rule that transforms the factorisation according to $\mathcal{G}$ to the factorisation according to $\mathcal{H}$ while ensuring the joint is the same. Separability-compatibility checks whether the prior separates according to multiple factorisations. If it does, ICM holds in multiple factorisations.
Theorem B.7.
Given two Bayesian causal models $(\mathcal{M}_{\mathcal{G}},\pi_{\mathcal{G}})$, and $(\mathcal{M}_{\mathcal{H}},\pi_{\mathcal{H}})$. If $\mathcal{G}$ and $\mathcal{H}$ are in the same Markov equivalence classes (MEC), and the Bayesian causal models are not separable compatible, they cannot be Bayesian distribution-equivalent. Further, assuming data is generated from one of the models, then the probability of error is less than $50\%$.
Proof.
The first statement follows directly from Dhir et al. (2024, Proposition 4.7). The probability of error in the two model case can be written as (Dhir et al., 2024)
$P(E)=\frac{1}{2}\mathopen{}\mathclose{{}\left(1-\text{TV}[P(\cdot|\mathcal{M}_{\mathcal{G}}),P(\cdot|\mathcal{M}_{\mathcal{H}})]}\right).$ (17)
If the two models are not Bayesian distribution-equivalent, then $\text{TV}[P(\cdot|\mathcal{M}_{\mathcal{G}}),P(\cdot|\mathcal{M}_{\mathcal{H}})]>0$. This implies that the probability of error is less than $\frac{1}{2}$. ∎
Theorem B.7 shows that if the ICM assumption only holds in the causal factorisation of the model (prior is only separable with respect to the causal factorisation), then there will always exist datasets that give different marginal likelihood values for the two Bayesian causal models. The above holds for any combination of Bayesian causal models with graphs in the same MEC. Hence, the total probability of error for the comparison within an MEC will be less than that of a graph randomly chosen from within the MEC (Friedman, 1996).
Appendix C CGP-CDE Details